OpenStudy (anonymous):

Explain how to write the equations of vertical lines, and why they are written this way

OpenStudy (amistre64):

use the standard form ....

OpenStudy (amistre64):

Ax + By = C if we were to try to solve this for y we get: y = -A/B x + C/B equating this to the usual y = mx + b form, we see that the slope is defined as -A/B but the slope of a vertical line is undefined, or rather, it is a division by 0, so B=0 substituting we find that Ax + 0y = C or simply, Ax = C as such x = C/A is the equation of a vertical line